OpenStudy (anonymous):

Find the distance for the following points A ( -2, -3) B (-7, -8). If necessary round your answer to the nearest tenth. (this math is killing me)

OpenStudy (anonymous):

yeah it is a pain

OpenStudy (anonymous):

Use the distance formula: sqrt((-2 - (-7))^2 + (-3 - (-8))^2) = sqrt( 25 + 25) = 5sqrt(2).

OpenStudy (anonymous):

1) 11 2) 7.1 3) 7 4) -7.1

OpenStudy (radar):

\[\sqrt{(-2--7)^{2}+(-3--8)^{2}}\]\[\sqrt{5^{2}+5^{2}}=\sqrt{50}=5\sqrt{2}=7.07\] Choice 3 is the closest.